OpenStudy (anonymous):

in how many ways can you draw a 6 or a face card from a deck of 52 cards

OpenStudy (anonymous):

yo, gurl. How many 6s are there? How many face cards are there?

OpenStudy (anonymous):

P(6) = 4/52 P(face card) = 12/52 P(6) or P(face card) = 4/52 + 12/52 = 16/52 = 4/13 = 0.307

OpenStudy (anonymous):

@fitzmaths...Please allow the person who asked the question to do *some* of the work. Otherwise, there's no incentive to learn.

OpenStudy (anonymous):

Fitz that kind of answer is worthless.

OpenStudy (anonymous):

i see what you mean. haha, my answer was about probability and the question asks about the number of way you can draw a card. I think a better answer would be simply 4 & 12

OpenStudy (anonymous):

@shandelman and @SmoothMath, i understand your point.

OpenStudy (anonymous):

The point is to just not solve the problem. Help them solve it on their own.

OpenStudy (anonymous):

@akita hope you can do more problems by yourself now

OpenStudy (anonymous):

It's cool. But you're right, the question was not about the probability, it was just about the number of ways. So, actually, the right answer has yet to be given.

OpenStudy (anonymous):

(Because the question was how many ways can you pick a 6 OR a face card? There's enough hints given that the answer is not too hard to figure out from what has been said).